1. Define pH and pKa. Use these concepts to predict the ionization state of any given amino acid or its side chain in a protein.
1
Expert's answer
2020-06-10T10:51:35-0400

pH is a measure of hydrogen ion concentration, a measure of the acidity or alkalinity of a solution. The pH scale ranges from 0 to 14. Aqueous solutions at 25°C with a pH less than 7 are acidic while those with a pH greater than 7 are basic or alkaline.

pH is an important quantity which reflects the chemical conditions of a solution. pH can control the availability of nutrients, biological functions, microbial activity and behaviour of chemicals. pH value of a food is a direct function of the free hydrogen ions present in that food. Acids prsent in foods relaese free hydrogen ions; the hydrogen ions give acid foods their distictive sour flavour. So, pH may be defined as a measure of free acidity. Stating precisely, pH is defined as the negative log of hydrogen ion concentartion. Range of pH extends from zero to fourteen. A pH value of 7 is neutral because pure water has a pH value of excatly 7. Values lower than 7 are acidic; values greater than 7 are basic or alkaline.

The pKa value is one method used to indicate strength of an acid. pKa is the negative log of acid dissociation constant or Ka value. Lower pKa value indicates a stronger acid. That is, lower value indicates that the acid dissociates more in water.

If pH or pKa values are known, one can solve for other value using an approximation called Henderson-Hasselbalch equation:

pH = pKa + log {[conjugate base]/[weak acid]}

pH = pKa + log{[A-]/[HA]

pH is the sum of the pKa value and the log of concentration of the conjugate base divided by concentration of weak acid.

At half the equivalence point: pH = pKa

pKa = -logKa

pKa values of amino acid side chains play a significant role in defining pH-dependent charcteristics of a protein. The pH-depnednece of the activity displayed by enzymes and the pH-dependence of protein stability, for instance, are properties which are determined by pKa values of amino side chains.

Whena protein folds. the titrable amino acids in the protein are transferred from a solution-like environment to an environment determined by the 3-dimensional structure of the protein. For example, in an unfolded peotein, an aspartic acid is in an environment which exposes the titrable side chain to water. When the protein folds, aspartic acid could find itswlf buried deep in protein interior with no exposure to solvent.

Moreover, in the folded protein, aspartic acid will be closer to other titrable groups in the protein and will also interact with permamnent charges like ions and dipoles in the protein. All these effects alter the pKa value of amino acid side chain and pKa calculation methods usually calculate the effect of protein environment on model pKa value of an amino acid side chain.

The efefcts of protein environment on the amino acid pKa value are divided into pH-independent effects and pH-independent effects. The pH-independent effects (desolvation, interactions with permanent charges and dipoles) are added to the model pKa value to give intrinsic pKa value. pH-independent effects cannot be added in same simple way and have to be accounted for using Boltzmann summation or Tanfors-Roxby iteration methods.

Seven out of the twenty amino acids contain readily ionizable side chain groups which means that at sepcific pH values, each side chain can participate in an acid-base raection in which it can exchange a hydrogen atom with some other biomolecule. Since these side chains can form ions, it implies that they can also participate in forming ionic bonds. On top of these seven amino acids with ionizable side chains, all of the amino acids contain an ionizable alpha amino group and an alpha carboxyl group. The pKa value of each side chain group determines pH value at which there will be equal concentartions of the acid and its conjugate base. The pKa value of alpha amino group is 3.1 whereas pKa value of alpha amino group is 8.0. pKa values of glutamic acid and aspartic acid are 4.1, pKa of cysteine is 6.0, pKa value of lysine is 8.3, pKa value of tyrosine is 10.9 and the pKa value of arginine is 12.5.

The isoelectronic point or isoionic point is the pH at which the amino acid does not migrate in an electric field.

This means it is the pH at which amino acid is neutral, i.e. the zwitterion form is dominant.

 There are three cases:

Neutral side chains: These amino acids are characterized by two pKas: pKa1 and pKa2 for carboxylic acid and the amine respectively.

The isoelectronic point will be halfway between, or average of, these two pKas , i.e. pI = 1/2 (pKa1 + pKa2). At very acidic pH (below pKa1), the amino acid will have an overall positive charge and at very basic pH ( below pKa1 ), the amino acis will have overall negative charge. For the simplest amino acid, gycine, below below pKa1 = 2.34 and below pKa1 = 9.6, pI = 5.97.

Acidic side chains: The pI will be at a lower pH because acidic side chain introduces an 'extra' negative charge. So, neutral form exists under more acidic conditiond when extra negative charge has been neutralized. For instance, for aspartic acid, neutral form is dominant between pH 1.88 and 3,65, pI is halfway between these two values, i.e. pI = 1/2 (pKa1 + below pKa3), so pI = 2.77.

Basic side chains: The pI will be at a higher pH because the basic side chain introduces an 'extra' positive charge. Thus, the neutral form exists under ore basic conditions when extra positive charge has been neutralized. For instance, in the case of histidine, the neutral form is dominant between pH 6.0 and 9.17, pI is halfway between these two values, i.e. pI = 1/2 (pKa2 + pKa3), so pI = 7.59.

Histidine is an alpha amino acid which is used in biosynthesis of proteins. It contains an alpha amino group (which is in protonated –NH3+ form) and an imidazole side chain (which is partially protonated), classifying it as a positively charged amino acid at physiological pH.

The conjugate acid (protonated form) of imidazole side chain in histidine has a pKa of nearly 6.0. So, at below a pH of 6, imidazole ring is protonated. Resulting imidazolium ring bears two NH bonds and has a positive charge which is equally distributed between both nitrogens and can be represented by two equally important resonance structures. Above pH = 6, one of the two protons is lost. Remaining proton of imidazole ring can reside on either nitrogen, giving rise to the N1-H or N3-H tautomers. The N3-H tautomer is protonated on #3 nitrogen, farther from amino acid backbone bearing the amino and carboxyl groups whereas N1-H tautomer is protonated on the nitrogen nearer to the backbone. The imidazole/imidazolium ring of histidine is aromatic at all pH values.